The current rate for producing 5 amp fuses at Neary Electric Co. is 250 per hour. A new machine has been purchased and installed that, according to the supplier, will increase the production rate. A sample of 10 randomly selected hours from last month revealed the mean hourly production on the new machine was 256 units, with a standard deviation of 6 per hour. At 5% significance level can we nearly conclude that the new machine is faster?


1
Expert's answer
2021-02-24T12:27:54-0500

We have that

"\mu = 250"

"n = 10"

"\bar x=256"

"s=6"

"\alpha=0.05"


"H_0:\mu = 250"

"H_a:\mu >250"

The hypothesis test is right-tailed.

Since the population standard deviation is unknown we use the t-test.

The critical value for 5% significance level and 9 df is 1.83

(degrees of freedom df = n – 1 = 10 – 1 = 9)

The critical region is t > 1.83

Test statistic:


"t=\frac{\bar x - \mu}{\frac{s}{\sqrt n}}=\frac{256- 250}{\frac{6}{\sqrt {10}}}=3.16"

Since 3.16 > 1.83 thus t falls in the rejection region we reject the null hypothesis.

At the 5% significance level the data do provide sufficient evidence to support the claim. We are 95% confident to conclude that that the new machine is faster.